# Predicting Subway Passenger Flows under Incident Situation with Causality

### v1

@@ Abstract
subway incident, passenger flow prediction, causal inference, significant test
journal: Elsevier
[…]
@@ 1 Introduction
The subway system is a critical component of the transportation network in metropolitan areas due to its environmental benefits, efficiency, and reliability. Understanding passenger flow patterns within these systems is essential for enhancing operational efficiency and user satisfaction. While much research has focused on typical operating conditions, there is a significant gap in addressing unusual situations, such as incidents. Analyzing and predicting how incidents impact passenger flows is vital for subway operators to implement proactive strategies. For instance, some studies have concentrated on emergency bus bridging and dispatching models during subway incidents [47, 39], considering the affected passenger numbers as inputs for their models. Consequently, we aim to fill this gap in our paper by proposing a method to interpret the influence of incident on OD (origination- destination) passenger flows and predict the passenger flows under incident situation.
@@
There are several papers [54, 26] that focus on predicting passenger flows in subway systems, treating incidents as input features. Since incidents are characteristics of the inputs, these models can predict passenger flows under both normal and incident conditions. They demonstrate improved predictive power and provide valuable insights. However, these models employ complex deep learning methods and end-to-end structures, which obscure the mechanisms by which incidents affect passenger flows. Furthermore, it can be challenging for these models to learn patterns related to incidents in depth, as the number of incidents in subway systems is typically small. Additionally, the imbalance between the volume of data in normal and incident situations may lead the model to prioritize normal conditions, hindering its ability to adequately learn the rules governing incident situations. Finally, the criteria for determining which OD flows are influenced by incidents have not been thoroughly discussed. Some methods [54] assume that the influenced OD flows occur during the incident period and the shortest paths of them include the incident stations. But this assumption is too simple and contradicts reality, for example, even though an incident ends, the effects of it may still persist. And we will elaborate it in the experiment section.
@@
To address these gaps, we propose a two-stage method for predicting passenger flow during incidents, as illustrated in Figure 1. First, a model trained with passenger flow data under normal conditions can be utilized to predict the OD flows that would occur in the absence of an incident. Subsequently, another model can be employed to predict the causal effects of the incident on passenger flows. These causal effects can then be incorporated into the normal OD flow predictions to obtain the final predictions. Besides, these causal effects can also be feed into emergency bus bridging and dispatching models [47, 39] to generate an appropriate bridging bus plan.
@@
This two-stage approach effectively addresses the aforementioned challenges to some extent. For instance, it offers a significant advantage in interpretability: it allows for a detailed analysis of the underlying mechanisms by which incidents impact passenger flows, as the incident model is separated. In contrast, an end-to-end model functions more like a “black box,” making it difficult to analyze the underlying mechanism. Furthermore, in a two-stage framework, high-quality data can be selectively utilized to calibrate the incident model, enhancing its ability to capture incident patterns. This process is more challenging for end-to-end models, as the incident data is often obscured within the normal data. Finally, our method provides flexibility, as each stage can be independently improved. For instance, if a more accurate normal OD prediction model becomes available in the future, it can be directly integrated into the first stage of our method.
@@
To identify the causal effects in historical data, the synthetic control framework is employed to estimate these effects [52]. Besides, to differentiate actual causal effects from noise, placebo tests are utilized to assess the statistical significance of these effects. Furthermore, to calibrate the causal effect prediction model, we examine which types of effects should be used to calibrate the model and propose that only significant effects should be considered. And it is essential to balance both the quantity and quality of data when determining the threshold of significant level. Finally, in the prediction phase, we analyze which types of ODs require correction by the causal effect prediction model and conclude that only significantly affected ODs should be adjusted. Both experimental and theoretical evidence supports the importance of selecting significantly affected ODs for model calibration and passenger flow prediction. We hope our research will provide valuable insights into the patterns of incident effects on passenger flows and offer practical tools for subway station staff to manage incidents effectively. We summarize our contributions as follows:
  • 1.
    We proposed a two-stage OD flow prediction model for subway system under incidents. The first stage is predicting the passenger number under normal situations and the second stage is combining this prediction with the causal effect of incident to obtain the final prediction.
  • 2.
    We used synthetic control method to identify the casual effect from history data and use placebo test to estimate its significance level.
  • 3.
@@ 2.1 OD passenger flow prediction models
In general, the main challenge of OD prediction is the high dimensionality [53]. For example, if there are 100 stations in a subway system, the number of ODs will be 10000, which could extremely challenging for traditional models.
@@
To address this, [5, 9] assumed the OD matrix is of low rank and can be reduced to some much smaller matrices. Therefore, some autoregression methods could be employed for these small matrices. And [53] assumed the proportion of passengers of one station to another station compared to the inflow number of this station is a constant, as a result, OD matrix can be directly obtain once the inflow prediction is available. [10] also used station-level embedding to obtain OD matrix and avoid the high dimension problem.
@@
In addition, deep learning method, which is more flexible, expressive and suitable for high dimensional data [15], was used in OD prediction models. For the aspect of time -axis modeling, [40] used LSTM to capture the time information in each time period. [8] employed discrete wavelet transform to capture the fluency information for OD flow in different time periods. And [46] proposed considering day and week periods is important. Besides attention mechanism [12] has been attempted too.
@@
As for spatial information, [6] utilized convolutional neuron network to capture patterns in history OD matrix. [11, 41, 21] attempted graph neuron networks. And [27] used wavelet transform to extract multi-resolution features in OD matrix. [19] introduced dual transformer to capture both temporary and spatial information and predict OD passenger flows. And [10] proposed a module to learn adaptive spatial relationships from data. In summary, even though there are a lot of papers addressing the challenge of OD passenger flow prediction, they often focus on normal conditions and ignore the effects of incidents.
2.2 Evaluating the effects of incident on subway systems
@@ 2.2.1 Network vulnerability
The majority of studies in this category utilize simulation-based methods or approaches grounded in complex network theory. These methodologies often focus on evaluating network redundancy and robustness. Complex network-based methods conceptualize the subway network as a graph, with stations as nodes and the connections between them as edges. By systematically removing nodes or edges, new graphs are generated, and researchers analyze changes in various coefficients that reflect the network’s structure or passenger travel convenience [7] This allows for the assessment of incident impacts in different regions and the overall resilience of the subway network.
@@
For instance,[49] analyzed Shanghai’s subway network by removing stations or intervals and evaluating changes in network efficiency, network size, average node centrality, and average edge centrality. This helped determine the vulnerability of specific intervals and assess the network’s overall robustness. Similarly, [48] used changes in network efficiency-related coefficients before and after node and edge deletions to evaluate the vulnerability of subway networks in several cities. Other notable studies include [30, 38]. [51] proposed various indicators to reflect network vulnerability, using concepts from causal inference and actual data to evaluate changes in these indicators under normal and incident scenarios.
@@
Additionally, some research incorporated network passenger flow into models to evaluate the effects of incidents on subway networks [25, 20]. Simulation-based methods were often employed to mimic passenger travel behavior under incident conditions, enabling the evaluation of passenger losses such as increased travel time and congestion costs [29]. And some work considered the cascading failure caused by one incident [28, 34] to evaluate the robustness of subway network, and found the some vulnerable stations or intervals. Furthermore, some articles proposed using simulations to quantify the spillover effects of incidents in specific intervals [45].
2.2.2 Incidents effect on passenger flows
@@
Some studies have utilized passenger flow prediction models to evaluate the effects of incidents. For example, [35] developed a prediction model for subway passenger flow in London based on normal conditions. This model can predict expected passenger flow in the absence of an incident, so this passenger flow can server as a counterfactual value to quantify the impact of subway incident. According to this thought, Other models designed for short-term passenger flow prediction, such as PVGCN [18] and MPSTN [14], could also be adapted to assess the effects of incidents on subway stations.
@@
Besides, [54] created an OD prediction model using smart card data, incorporating incidents as an independent variable. This model can calculate OD values under both incident and non-incident conditions, with the difference indicating the impact of subway incidents on passenger flows. Additional studies have developed specific prediction models for passenger flow during incidents, comparing these results with normal conditions to gauge the impact [42, 43, 24]. Simulation-based approaches have also been used to estimate passenger flow at stations during incidents, thereby evaluating the impact at various stations [44, 36].
Furthermore, [52] utilized real operational data and causal inference methods to assess the effects of subway system disruptions on station passenger flows, highlighting the spillover effects of some disruptions.
@@
Several studies have also analyzed and predicted individual passenger behaviors during incidents, which can be extended to all passengers and evaluate the change of passenger flow number due to incidents. Some of these works used survey data to understand passenger choices during incidents [32, 22]. Other research leveraged transport card data in consecutive days to identify travel behaviors under incident conditions, and compared these to normal conditions and established predictive models for passenger behavior changes during incidents [37]. Additionally, [23] predicted how many passengers would alter their travel behavior and how these passengers would change their behaviors during incidents.
@@ 2.3 Summary
In summary, very few studies address passenger flow prediction for subway systems under incident conditions. And these existing works primarily apply deep learning models, often with simplistic assumptions about which passenger flows are impacted by subway incidents. These studies do not delve into how incidents affect passenger flows in detail, and the interpretability of their models is generally limited.
Moreover, most existing literature about the impact of incidents on subway systems focuses on network vulnerability, often relying on simulations or complex network methods that may not fully reflect real-world subway operations. Besides, their research questions are pretty different compared with the questions proposed in this work. Some studies evaluate the impact of incidents on passenger flow using predictive models, comparing predicted normal conditions with actual passenger flow during incidents. However, these studies lack analysis from causality. Besides, they can only calculate these differences after incidents ended.
@@ 3.1 Causal effect estimation
We have observed that some research focuses on identifying the causal effects of highway incidents on traffic state [16], and they used double robust or inverse propensity weighting methods to address selection biases, i.e., whether an incident happens is related to the influence factors of traffic state. However, incidents in subway systems differ significantly from those on highways, with one of the most notable differences being the number of incidents. The frequency of incidents in subway systems is considerably lower than that in highways. Furthermore, for a specific OD, not all incidents will have an impact on it. Consequently, the number of incidents affecting a specific OD may be even smaller. Therefore, the adjustment of selection biases may be unnecessary. But we require a method to address the issue of having only a limited number of experimental samples.
We use the synthetic control method [2] to calculate counterfactual values and obtain causal effects. This method is widely applied in empirical research within economics and social sciences [1] when only a few samples are intervened. Following the algorithm proposed in [2], we design our approach to estimate the causal effects. But at first, we will give some notations.
@@
$\begin{array}[]{cccccccc}x_{1,1}^{0}&x_{1,2}^{0}&\cdots&x_{1,k-1}^{0}&x_{1,k}^{0}&x_{l,k+1}^{0}&\cdots&x_{1,m}^{0}\\ x_{2,1}^{0}&x_{2,2}^{0}&\cdots&x_{2,k-1}^{0}&x_{2,k}^{0}&x_{2,k+1}^{0}&\cdots&x_{2,m}^{0}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ x_{l-1,1}^{0}&x_{l-1,2}^{0}&\cdots&x_{l-1,k-1}^{0}&x_{l-1,k}^{0}&x_{l-1,k+1}^{0}&\cdots&x_{l-1,m}^{0}\\ x_{l,1}^{0}&x_{l,2}^{0}&\cdots&x_{l,k-1}^{0}&\bm{x_{l,k}^{1}}&\bm{x_{l,k+1}^{1}}&\cdots&\bm{x_{l,m}^{1}}\\ x_{l+1,1}^{0}&x_{l+1,2}^{0}&\cdots&x_{l+1,k-1}^{0}&x_{l+1,k}^{0}&x_{l+1,k+1}^{0}&\cdots&x_{l+1,m}^{0}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ x_{n,1}^{0}&x_{n,2}^{0}&\cdots&x_{n,k-1}^{0}&x_{n,k}^{0}&x_{n,k+1}^{0}&\cdots&x_{n,m}^{0}\\ \end{array}$ (1)
@@
For convenience, we use W to signify the vector composed of all $w_{i}$’s: $W=(w_{1},…w_{l-1},w_{l+1},…,w_{n})\in\mathbb{R}^{n-1}$. The challenge then becomes determining $w_{i}$. Recall that a $d-$dimensional vector $a_{i,j}$ is defined to represent the influence factors of $x_{i,j}$. Therefore, the target of determining $W$ is to make the variable $a$ of the synthetic sample more similar to the $a$ in day $l$. Let $A$ be the matrix of all $a_{i,j}$’s where $i\in\{1,2,…,l-1,l+1,…,n\}$. Thus, $A\in\mathbb{R}^{n-1,d}$. Then $W$ can be determined as:
@@
But the importance of each dimension in $a$ may not be equal. As a result, [1] propose to use $V-$norm instead of $l_{2}$ norm and we follow this setting in our method. So $W$ is determined by:
@@
where $V$ is a $d\times d$ positive diagonal matrix. Equation 4 means the influence factors of synthetic counterfactual data, i.e., $WA$, are similar to the influence factors of actual counterfactual data $a_{l,k}$. The matrix $V$ represents the importance of each dimension in a. According to [1], $V$ can be chosen to minimize the prediction error in previous $t$ time intervals, as the following Equation 5 shows:
@@
where $X_{0}$ is a $t-$dimensional vector, $X_{0}=(x_{l,k-t}^{0},x_{l,k-t+1}^{0},…,x_{l,k-1}^{0})$, and $X$ is a matrix with shape $t\times(n-1)$, the $i$-th column of $X$ representing the passenger flow number in time intervals $k-t$ to $k-1$ of day $i$. $W(V)$ is the function mapping $V$ to $W$, as defined by Equation 4. For more details, interested readers can refer to [1]. Finally, the causal effect can be estimated as $x_{l,k}^{1}-\hat{x}_{l,k}^{0}$.
Some readers may question that why we do not choose to fit a model form $a$ to $x^{0}$ directly but use such a complicated process. This is because directly fitting model could meet the problem of extrapolation and results in inferior performance, this phenomenon has been discussed by [3] and interested readers can refer to it.

### v2

@@ Abstract
subway incident, passenger flow prediction, causal inference, synthetic control method
@@ 1 Introduction
The subway system is a critical component of the transportation network in metropolitan areas due to its environmental benefits, efficiency, and reliability. Understanding passenger flow patterns within these systems is important for enhancing operational efficiency and user satisfaction. While much research has focused on typical operating conditions, there is a significant gap in addressing unusual situations, such as incidents. Analyzing and predicting how incidents impact passenger flows is vital for subway operators to implement proactive strategies. For instance, some studies about emergency bus bridging and dispatching models during subway incidents use the affected passenger numbers as inputs for their models [Zhang2023Resi, Tan2024EnhancingUM]. Consequently, we aim to fill this gap in our paper by proposing a method to interpret the influence of incident on OD (origination- destination) passenger flows and predict the passenger flows under incident situation.
@@
There are several papers [Zou2024RealtimePO, NOURSALEHI2018277] that focus on predicting passenger flows in subway systems, treating incidents as input features. Since incidents are characteristics of the inputs, these models can predict passenger flows under both normal and incident conditions. They demonstrate improved predictive power and provide valuable insights. However, these models employ complex deep learning methods and end-to-end structures, which might obscure the mechanisms by which incidents affect passenger flows. Furthermore, it can be challenging for these models to learn patterns related to incidents in depth, as the number of incidents in subway systems is typically small. Additionally, the imbalance between the volume of data in normal and incident situations may lead the model to prioritize normal conditions, hindering its ability to adequately learn the rules governing incident situations. Finally, the criteria for determining which OD flows are influenced by incidents have not been thoroughly discussed. Some methods [Zou2024RealtimePO] assume that the influenced OD flows only occur during the incident period. But this assumption is too simple and contradicts reality, for example, even though an incident ends, the effects of it may still persist. And we will elaborate it in the experiment section.
@@
To address these gaps, we propose a two-stage method for predicting passenger flow during incidents, as illustrated in Figure 1. First, a model trained with passenger flow data under normal conditions can be utilized to predict the OD flows that would occur in the absence of an incident. Subsequently, another model can be employed to predict the causal effects of the incident on passenger flows. These causal effects can then be incorporated into the normal OD flow predictions to obtain the final predictions. Besides, these causal effects can also be fed into emergency bus bridging and dispatching models [Zhang2023Resi, Tan2024EnhancingUM] to generate an appropriate bridging bus plan.
@@
This two-stage approach effectively addresses the aforementioned challenges to some extent. For instance, this approach offers significantly improved interpretability. By separating the incident model, we enable detailed analysis of the underlying mechanisms by which incidents impact passenger flows. In contrast, an end-to-end model functions more like a “black box,” making it difficult to analyze the underlying mechanism. Furthermore, in a two-stage framework, high-quality data can be selectively utilized to calibrate the incident model, enhancing its ability to capture incident patterns. This process is more challenging for end-to-end models, as the incident data is often obscured within the normal data. Finally, our method provides flexibility, as each stage can be independently improved. For instance, if a more accurate normal OD prediction model becomes available in the future, it can be directly integrated into the first stage of our method.
@@
To identify the causal effects in historical data, the synthetic control framework is employed to estimate these effects [zhang2023causalinferencedisruptionmanagement]. Besides, to differentiate actual causal effects from noise, placebo tests are utilized to assess the statistical significance of these effects. Furthermore, to calibrate the causal effect prediction model, we examine which types of effects should be used to calibrate the model and propose that only significant effects should be considered. And it is essential to balance both the quantity and quality of data when determining the threshold of significant level. Finally, in the prediction phase, we analyze which types of ODs require correction by the causal effect prediction model and conclude that only significantly affected ODs should be adjusted. Both experimental and theoretical evidence supports the importance of selecting significantly affected ODs for model calibration and passenger flow prediction. We hope our research will provide valuable insights into the patterns of incident effects on passenger flows and offer practical tools for subway station staff to manage incidents effectively. We summarize our contributions as follows:
  • We proposed a two-stage OD flow prediction model for subway system during incidents. The first stage predicts the passenger number under normal situations and the second stage combines this prediction with the causal effect of incident to obtain the final prediction.
  • We used synthetic control method to identify the causal effect from historical data and used placebo test to estimate its significance level.
@@ 2.1 OD passenger flow prediction models
In general, the main challenge of OD prediction is the high dimensionality [Two-Stage2024]. For example, if there are 100 stations in a subway system, the number of ODs will be 10000, which is extremely challenging for traditional models.
@@
To address this, [Cheng2021RealTimeFO, Gong2020OnlineSC] assumed that the OD matrix is of low rank and can be reduced to some much smaller matrices. Therefore, some autoregression methods could be employed for these small matrices. [Two-Stage2024] assumed that the proportion of passengers traveling from one station to another, relative to the inflow number of the origin station, remains constant. As a result, the OD matrix can be directly obtained once the inflow prediction is available. [Han2022ContinuousTimeAM] also used station-level embedding to estimate OD matrix and avoid the high dimension problem.
@@
In addition, deep learning method, which is more flexible, expressive and suitable for high dimensional data [lecun_deep_2015], was used in OD prediction models. For the aspect of time -axis modeling, [Toque2016ForecastingDP] used LSTM to capture temporal patterns. [HybridTITS2019] employed discrete wavelet transform to capture the fluency information for OD flow in different time periods. And [Ye2022CompletionAA] proposed considering day and week periods is important. Besides, attention mechanism [HUANG2023ODformer] have been attempted too.
@@
As for spatial information, [Chu2020DeepMC] utilized convolutional neuron network to capture patterns in history OD matrix. [Stochastic2020, Wang2019OriginDestinationMP, Makhdomi2023GNNbasedPR] attempted graph neuron networks. And [Noursalehi2021DynamicOP] used wavelet transform to extract multi-resolution features in OD matrix. [Liu2021OnlineMO] introduced dual transformer to capture both temporary and spatial information and predict OD passenger flows. And [Han2022ContinuousTimeAM] proposed a module to learn adaptive spatial relationships from data. In summary, despite a lot of research addressing the challenge of OD passenger flow prediction, these methods often focus on normal conditions and fail to account for the effects of incidents.
2.2 Causality method for traffic prediction and incident effects
Previous research has applied causality methods to predict traffic flows or traffic demand. For example, [li2025causal] proposed a spatiotemporal causal intervention-based large language model approach to eliminate spurious spatial correlations and mitigate hallucination issues. [li2015robust] employs Granger causality theory to select the most relevant parts from massive traffic data to build concise prediction models. [khaled2024graph] uses transfer entropy to describe the causal relationship between each pair of nodes, constructs a causal adjacency matrix, and feeds it into a spatiotemporal prediction model for traffic forecasting. [xu2024traffic] proposed a Multisource Causal Interpretation Graph based reasoning model to infer traffic state data for road sections without detectors. These works are highly valuable. However, their focus is conducting traffic prediction under normal situations, which differs from our research topic.
Besides, there are also some works focusing on analyzing the impact of highway incident using causal inference methods. For instance, [li2024inferring] analyzed the causal effects of highway incidents on subsequent traffic conditions. Factors like weather and rush hour periods are regarded as confounding variables that simultaneously influence both the probability of incident occurrence and the resulting effects. Doubly robust learning methods are employed to eliminate the influence of these confounding factors. Similar ideas can be found in [zhang2022inferring, li2016quantifying, cao2021quantification]. However, highway incidents are relatively frequent, and the datasets used in these works include thousands of incidents, enabling the calibration of a propensity score model to predict incident probability based on confounders. In contrast, subway systems are generally more stable, and our dataset contains only six incidents with significant impact. Therefore, calibrating a reliable propensity score model is highly challenging. Secondly, most subway incidents in our dataset involve faults like communication system failures or station electromechanical facility malfunctions. These may not exhibit strong confounding effects – meaning there are likely no variables that simultaneously significantly influence both the probability of an incident occurring and the impact of that incident. Therefore, even without a propensity score model might be acceptable. However, we need to address the challenge of data sparsity, where the number of treated samples (incident cases) is very small. Consequently, in this situation, we opted to use the Synthetic Control Method, which is specifically designed for scenarios with a limited number of intervention cases.
Furthermore, [zhang2023causalinferencedisruptionmanagement] utilized real operational data and causal inference methods to assess the effects of subway system disruptions on station passenger flows, highlighting the spillover effects of some disruptions. However, it focuses on station-level passenger flow and only identifies the causal effects of incidents, lacking a predictive module for incident impacts. This approach is more suitable for post-event analysis rather than real-time passenger flow prediction.
2.3 Evaluating the effects of incidents on subway systems
@@ 2.3.1 Network vulnerability
The majority of studies in this category utilize simulation-based methods or approaches grounded in complex network theory. These methodologies often focus on evaluating network redundancy and robustness. Complex network-based methods conceptualize the subway network as a graph, with stations as nodes and the connections between them as edges. By systematically removing nodes or edges, new graphs are generated, and researchers analyze changes in various coefficients that reflect the network’s structure or passenger travel convenience [Derrible2010TheCA] This allows for the assessment of incident impacts in different regions and the overall resilience of the subway network.
@@
For instance,[Zhang2011NetworkedAO] analyzed Shanghai’s subway network by removing stations or intervals and evaluating changes in network efficiency, network size, average node centrality, and average edge centrality. This helped determine the vulnerability of specific intervals and assess the network’s overall robustness. Similarly, [Zhang2018ComparisonAO] used changes in network efficiency-related coefficients before and after node and edge deletions to evaluate the vulnerability of subway networks in several cities. Other notable studies include [Qi2022ResilienceAO, Sun2018VulnerabilityAO]. [Zhang2020ACI] proposed various indicators to reflect network vulnerability, using concepts from causal inference and actual data to evaluate changes in these indicators under normal and incident scenarios.
@@
Additionally, some research incorporated network passenger flow into models to evaluate the effects of incidents on subway networks [Nian2019EvaluatingTA, Liu2024AssessingMN]. Simulation-based methods were often employed to mimic passenger travel behavior under incident conditions, enabling the evaluation of passenger losses such as increased travel time and congestion costs [paulsen_modelling_2018]. And some work considered the cascading failure caused by one incident [Pan2024OnTD, Shen2019CascadingFA] to evaluate the robustness of subway network, and found the some vulnerable stations or intervals. Furthermore, some articles proposed using simulations to quantify the spillover effects of incidents in specific intervals [Yap2021QuantificationAC].
2.3.2 Incident effects on passenger flows
@@
Some studies have utilized passenger flow prediction models to evaluate the effects of incidents. For example, [Silva2015PredictingTV] developed a prediction model for subway passenger flow in London under normal conditions. This model can predict expected passenger flow in the absence of an incident, so this passenger flow can serve as a counterfactual value to quantify the impact of subway incident. Based on this concept, other models designed for short-term passenger flow prediction, such as PVGCN [Liu2020PhysicalVirtualCM] and MPSTN [huang_leveraging_2024], could also be adapted to assess the effects of incidents on subway stations.
@@
Besides, [Zou2024RealtimePO] created an OD prediction model using smart card data, incorporating incident status as an independent variable. This model can calculate OD values under both incident and non-incident conditions, with difference indicating the impact of subway incidents on passenger flows. Other studies have developed specific prediction models for passenger flow during incidents, comparing these outputs with predictions under normal conditions to measure the impact [Xu2023ShorttermPF, Xue2021ForecastingTS, Ni2017ForecastingTS]. Simulation-based approaches have also been used to estimate passenger flow during incidents, thereby evaluating the impact at various stations [Yap2020PredictingDA, Su2023SimulationBasedMF].
@@
Several studies have also analyzed and predicted individual passenger behaviors during incidents, which can be extended to all passengers and evaluate the change of passenger flow number due to incidents. Some of these works used survey data to understand passenger choices during incidents [Rubin2005PsychologicalAB, Masud2019TravelBR]. Other research leveraged transport card data in consecutive days to identify travel behaviors under incident conditions, and compared these to normal conditions and established predictive models for passenger behavior changes during incidents [Sun2016EstimatingTI]. Additionally, [Mo2022InferringPR] predicted how many passengers would alter their travel behavior and how these passengers would change their behaviors during incidents.
@@ 2.4 Summary
In summary, existing works about passenger flow prediction under incident situation primarily apply deep learning models, often with simplistic assumptions about which passenger flows are impacted by subway incidents. These studies do not delve into how incidents affect passenger flows in detail, and the interpretability of their models is generally limited. Besides, other methods using passenger choice models to infer subway passenger flow under incident conditions typically rely on socioeconomic attributes, which may be unavailable during real-time subway operations. Studies using simulation methods for passenger flow evolution might impose overly simplistic or unrealistic rules on passenger behavior. Therefore, our paper adopts real passenger flow data, which could be complement to them.
Literature about the impact of incidents on subway systems focuses on network vulnerability, often relying on simulations or complex network methods that may not fully reflect real-world subway operations. Besides, their research questions are different compared with the questions proposed in this work. Some studies evaluate the impact of incidents on passenger flow using predictive models, comparing predicted normal conditions with actual passenger flow during incidents. However, these studies lack causality analysis. Additionally, they can only calculate these differences after incidents ended.
@@ 3.1 Causal effect estimation
We use the synthetic control method [Abadie2010SyntheticCM] to calculate counterfactual values and obtain causal effects. This methods have been used in transportation section and many researchers used them to evaluate policy effects. For example, [yang2022assessing, dai2021improving] assess the effects of subway price policy on passenger flows and [wallimann2023price] evaluate the effects of new railway lines on congestion. Following the algorithm proposed in [Abadie2010SyntheticCM], we designed our approach to estimate the causal effects. At first, we will give some notations.
@@
[…]
@@
For convenience, we use W to signify the vector composed of all $w_{i}$’s: $W=(w_{1},…w_{l-1},w_{l+1},…,w_{n})\in R^{n-1}$. The challenge then becomes determining $w_{i}$. Recall that a $d$-dimensional vector $a_{i,j}$ is defined to represent the influence factors of $x_{i,j}$. Therefore, the target of determining $W$ is to make the variable $a$ of the synthetic sample more similar to the $a$ in day $l$. Let $A$ be the matrix of all $a_{i,j}$’s where $i\in\{1,2,…,l-1,l+1,…,n\}$. Thus, $A\in R^{n-1,d}$. Then $W$ can be determined as:
@@
But the importance of each dimension in $a$ may not be equal. As a result, [Abadie2021UsingSC] propose to use $V$-norm instead of $l_{2}$ norm and we follow this setting in our method. So $W$ is determined by:
@@
where $V$ is a $d\times d$ positive diagonal matrix. Equation 4 means the influence factors of synthetic counterfactual data, i.e., $WA$, are similar to the influence factors of actual counterfactual data $a_{l,k}$. The matrix $V$ represents the importance of each dimension in a. According to [Abadie2021UsingSC], $V$ can be chosen to minimize the prediction error in previous $t$ time intervals, as the following Equation 5 shows:
@@
where $X_{0}$ is a $t$-dimensional vector, $X_{0}=(x_{l,k-t}^{0},x_{l,k-t+1}^{0},…,x_{l,k-1}^{0})$, and $X$ is a matrix with shape $t\times(n-1)$, the $i$-th column of $X$ representing the passenger flow number in time intervals ktk-t to k1k-1 of day $i$. $W(V)$ is the function mapping $V$ to $W$, as defined by Equation 4. For more details, interested readers can refer to [Abadie2021UsingSC]. Finally, the causal effect can be estimated as $x_{l,k}^{1}-\hat{x}_{l,k}^{0}$.
Some readers may question that why we do not choose to fit a model form $a$ to $x^{0}$ directly but use such a complicated process. This is because directly fitting model could meet the problem of extrapolation and results in inferior performance, this phenomenon has been discussed by [Abadie2014ComparativePA] and interested readers can refer to it.